OpenStudy (anonymous):

Write an equation in point slope form of the line that passes through the given points j(4, 1) with slope -4

OpenStudy (haichi):

y-1=-4(x-4) medal and fan plz?

OpenStudy (haichi):

Slope point form=y-y1=m(x-x1)

OpenStudy (anonymous):

I did c: Thanks!

OpenStudy (haichi):

did u pass? and np